-2x^3-12x^2+30x+200=0
Answer by fractalier(6550)

This factors out as, after dividing by -2...
-2x^3-12x^2+30x+200=0
x^3 + 6x^2 - 15x - 100 = 0
(x+5)(x+5)(x-4) = 0
and
x = -5 or x = 4